OpenStudy (anonymous):

What is the equation of a line that passes through (7, 8) and has a slope of -3? y=−3x+29 y=3x+13 y=1/3x−29 y=−1/3x−13

OpenStudy (texaschic101):

y = mx + b slope(m) = -3 (7,8)...x = 7 and y = 8 now we sub and solve for b 8 = -3(7) + b 8 = -21 + b 21 + 8 = b 29 = b so our equation is : y = -3x + 29